# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ 3.1 Equivalence between Gaussian measures
If $w_{i}\in\operatorname{dom}{\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}}=\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ for some $i$, then the statement of item (i) implies $\mathcal{C}_{2}\mathcal{C}_{1}^{-1/2}w_{i}=(1+\lambda_{i})\mathcal{C}_{1}^{1/2}w_{i}$. In other words, $\mathcal{C}_{2}(\mathcal{C}_{1}^{-1/2}w_{i})=\mathcal{C}_{1}(\mathcal{C}_{1}^{-1/2}w_{i})$, showing that $(1+\lambda_{i},\mathcal{C}_{1}^{-1/2}w_{i})$ is a generalised eigenpair of the generalised operator pencil $(\mathcal{C}_{\textup{pos}},\mathcal{C}_{\textup{pr}})$. Furthermore, by the relation $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}w_{i}$, it follows that $v_{i}\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$, and the relation item (ii) implies $\mathcal{C}_{1}^{-1}\mathcal{C}_{2}^{1/2}v_{i}=(1+\lambda_{i})\mathcal{C}_{2}^{-1}\mathcal{C}_{2}^{1/2}v_{i}$ showing that $(1+\lambda_{i},\mathcal{C}_{2}^{1/2}v_{i})=(1+\lambda_{i},\sqrt{1+\lambda_{i}}\mathcal{C}_{1}^{1/2}w_{i})$ is a generalised eigenpair of $(\mathcal{C}_{1}^{-1},\mathcal{C}_{2}^{-1})$. Thus, in the case $(w_{i})_{i}$ and $(v_{i})_{i}$ lie in the Cameron–Martin space $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$, items (i), (ii), (iii) and (iv) in Lemma 3.3 can be interpreted as statements on operator pencils. The statements in Lemma 3.3 are, however, more general.
@@
Lemma 3.6.
There exists an increasing sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ consisting of exactly $\operatorname{rank}\left({H}\right)$ nonzero elements and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for every $i\in\mathbb{N}$, and
@@
In particular, it holds that
$\frac{\textup{Var}_{X\sim\mu_{\textup{pos}}}(\langle X,\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\rangle)}{\textup{Var}_{X\sim\mu_{\textup{pr}}}(\langle X,\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}\rangle)}=1+\lambda_{i}=\frac{1}{1+\tfrac{-\lambda_{i}}{1+\lambda_{i}}},\quad\forall i\in\mathbb{N}.$ (11)
Remark 3.7 (Variance reduction).
The significance of (11) is that the posterior variance along the span of $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ is smaller than the prior variance along the same subspace by a factor of $1+\tfrac{-\lambda_{i}}{1+\lambda_{i}}$, for $i\in\mathbb{N}$. This was observed in the finite-dimensional case by [30, Equation (3.4)]. Thus, Lemma 3.6 implies that finite-dimensional data can only inform finitely many directions in parameter space, in the sense that posterior variance is reduced relative to prior variance only over a finite-dimensional subspace. The directions $(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})_{i\leq\operatorname{rank}\left({H}\right)}$ are orthogonal with respect to the $\mathcal{C}_{\textup{pr}}$-weighted inner product $\langle h_{1},h_{2}\rangle_{\mathcal{C}_{\textup{pr}}}\coloneqq\langle\mathcal{C}_{\textup{pr}}h_{1},h_{2}\rangle$, and not the unweighted inner product of $\mathcal{H}$. If, for some $r<\operatorname{rank}\left({H}\right)$, there exists an $r$-dimensional subspace and the variance reduction on the complement of this subspace is sufficiently small, then this $r$-dimensional subspace is also called the ‘likelihood-informed-subspace’ in literature, see e.g. [8, 9, 10].
In Lemma 3.6, $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$, hence by Remark 3.5, the equations (10a) and (10b) can be interpreted as statements on operator pencils. More specifically, for any $i$, (10a) states that $(\frac{-\lambda_{i}}{1+\lambda_{i}},\mathcal{C}_{\textup{pr}}^{1/2}w_{i})$ is a generalised eigenpair of $(H,\mathcal{C}_{\textup{pr}}^{-1})$ and (10b) states that $(-\lambda_{i},\mathcal{C}_{\textup{pos}}^{1/2}v_{i})=(-\lambda_{i},\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})$ is a generalised eigenpair of $(H,\mathcal{C}_{\textup{pos}}^{-1})$. Furthermore, (10c) can be interpreted as a statement on the operator pencil $(\mathcal{C}_{\textup{pos}},\mathcal{C}_{\textup{pr}})$. The prior-preconditioned Hessian $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}$ has been found to be the central object of study in the reduction of finite-dimensional linear Gaussian inverse problems, see [30, 8]. We observe that this operator is directly related to $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ via the equivalent characterisations given by Lemma 3.3 Items (i), (ii), (iii) and (iv) and hence to the function $R(\cdot\|\cdot)$ which quantifies the similarity of Gaussian measures by Theorem 3.1(iii).
@@ 3.2 Divergences between equivalent Gaussian measures
To measure the quality of an approximation $\tilde{\mathcal{C}}\in\mathcal{E}$ of $\mathcal{C}_{\textup{pos}}$ and $\tilde{m}_{\textup{pos}}$ to $m_{\textup{pos}}$, we shall use the Rényi divergences of order $\rho\in(0,1)$ and the forward an reverse Kullback-Leibler divergence (KL-divergence). The KL-divergence from a measure $\nu_{1}$ to a measure $\nu_{2}$ is defined as
@@
If $\nu_{1}$ is an approximation of $\nu_{2}$, then the forward KL-divergence of the approximation is $D_{\textup{KL}}(\nu_{2}\|\nu_{1})$ and the reverse KL-divergence is $D_{\textup{KL}}(\nu_{1}\|\nu_{2})$. The Rényi divergence of order $\rho\in(0,1)$ is defined by, c.f. [22, Equation (130)],
$D_{\textup{Ren},\rho}(\nu_{2}\|\nu_{1})=-\frac{1}{\rho(1-\rho)}\log{\int_{\mathcal{H}}\left(\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}\right)^{\rho}\operatorname{d}\!{\nu}_{1}}$
In the Gaussian case, an explicit representation of these divergences holds, as shown in [22]. For this, we need a generalisation of the determinant to infinite-dimensional Hilbert spaces. Because in infinite dimensions the eigenvalues of a compact operator accumulate at 0, direct extension of the finite-dimensional definition of the determinant as the product of the eigenvalues to infinite dimensions will result in the determinant function being equal to the constant 00. A generalisation of the concept of the determinant for trace-class and Hilbert–Schmidt operators is given by the Fredholm determinant and Hilbert–Carleman determinant respectively. These are defined on respectively trace-classs and Hilbert–Schmidt perturbations of the identity, and are indicated with $\det(I+A)$, $A\in L_{1}(\mathcal{H})$, and respectively $\det_{2}{(I+A)}$, $A\in L_{2}(\mathcal{H})$. We refer to [27, Theorem 3.2, Theorem 6.2] or [28, Lemma 3.3, Theorem 9.2]. For $A\in L_{1}(\mathcal{H})$, we have $\det_{2}{(I+A)}=\det(I+A)\exp(-\operatorname{tr}\,\left({A}\right))$, and for $A\in L_{2}(\mathcal{H})$ the determinants are related via $\det_{2}(I+A)=\det(I+(I+A)\exp(-A)).$ By [27, Theorem 4.2, Theorem 6.2] or [28, Theorem 3.7, Theorem 9.2] for each $\mu\in\mathbb{R}$, we have the expression
@@
where $(\lambda_{i})_{i}$ denotes the eigenvalue sequence of $A$. In the case that $\dim{\mathcal{H}}<\infty$ we note that $A-I\in L_{1}(\mathcal{H})$ and $\det(A)=\det(I+(A-I))=\prod_{i}\lambda_{i}$ and thus $\det(\cdot)$ indeed extends the finite-dimensional definition of the determinant. We can now formulate the explicit expression of the KL and Rényi divergences for equivalent Gaussian measures.
@@
Let $m_{1},m_{2}\in\mathcal{H}$ and $\mathcal{C}_{1},\mathcal{C}_{2}\in L_{2}(\mathcal{H})_{\mathbb{R}}$ be nonnegative. If $m_{1}-m_{2}\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ and if $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ satisfies property E, then

### v3

@@ 3.1 Equivalence between Gaussian measures
If $w_{i}\in\operatorname{dom}{\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}}=\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ for some $i$, then the statement of item (i) implies $\mathcal{C}_{2}\mathcal{C}_{1}^{-1/2}w_{i}=(1+\lambda_{i})\mathcal{C}_{1}^{1/2}w_{i}$. In other words, $\mathcal{C}_{2}(\mathcal{C}_{1}^{-1/2}w_{i})=(1+\lambda_{i})\mathcal{C}_{1}(\mathcal{C}_{1}^{-1/2}w_{i})$, showing that $(1+\lambda_{i},\mathcal{C}_{1}^{-1/2}w_{i})$ is a generalised eigenpair of the generalised operator pencil $(\mathcal{C}_{2},\mathcal{C}_{1})$. Furthermore, if $(v_{i})_{i}$ lies in the dense subspace $\operatorname{dom}{\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1}\mathcal{C}_{2}^{1/2}}=\operatorname{dom}{\mathcal{C}_{1}^{-1}\mathcal{C}_{2}^{1/2}}$, then for any $i$ we have $\mathcal{C}_{2}^{1/2}v_{i}\in\operatorname{dom}{\mathcal{C}_{1}^{-1}}$. The relation in item (ii) shows that $\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1}\mathcal{C}_{2}^{1/2}v_{i}=(1+\lambda_{i})v_{i}$, so that $v_{i}\in\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$. Hence, $\mathcal{C}_{2}^{1/2}v_{i}\in\operatorname{dom}{\mathcal{C}_{1}^{-1}}\cap\operatorname{dom}{\mathcal{C}_{2}^{-1}}$. The previous relation implies $\mathcal{C}_{1}^{-1}\mathcal{C}_{2}^{1/2}v_{i}=(1+\lambda_{i})\mathcal{C}_{2}^{-1}\mathcal{C}_{2}^{1/2}v_{i}$, showing that $(1+\lambda_{i},\mathcal{C}_{2}^{1/2}v_{i})=(1+\lambda_{i},\sqrt{1+\lambda_{i}}\mathcal{C}_{1}^{1/2}w_{i})$ is a generalised eigenpair of $(\mathcal{C}_{1}^{-1},\mathcal{C}_{2}^{-1})$. Thus, in the case $(w_{i})_{i}$ and $(v_{i})_{i}$ lie in a dense set of $\mathcal{H}$, items (i), (ii), (iii) and (iv) in Lemma 3.4 can be interpreted as statements on operator pencils. The statements in Lemma 3.4 do not assume that $(w_{i})_{i}$ and $(v_{i})_{i}$ are contained in the particular dense subspaces of $\mathcal{H}$ on which the leftmost operators are defined. Therefore, these statements generalise the interpretation of a generalised eigenpencil given above.
@@
Proposition 3.7.
There exists a nondecreasing sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ consisting of exactly $\operatorname{rank}\left({H}\right)$ nonzero elements and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for every $i\in\mathbb{N}$, and
$R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ $=\sum_{i}\lambda_{i}w_{i}\otimes w_{i},$
@@
In Proposition 3.7, $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$, so that $v_{i}\in\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}}$ and $w_{i}\in\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}^{1/2}}$, because $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$. The equations (10a) and (10b) can be interpreted as statements on operator pencils by Remark 3.6. More specifically, (10a) states that $(\frac{-\lambda_{i}}{1+\lambda_{i}},\mathcal{C}_{\textup{pr}}^{1/2}w_{i})$ is a generalised eigenpair of $(H,\mathcal{C}_{\textup{pr}}^{-1})$ and (10b) states that $(-\lambda_{i},\mathcal{C}_{\textup{pos}}^{1/2}v_{i})=(-\lambda_{i},\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})$ is a generalised eigenpair of $(H,\mathcal{C}_{\textup{pos}}^{-1})$, for any $i$. Furthermore, (10c) can be interpreted as a statement on the operator pencils $(\mathcal{C}_{\textup{pos}},\mathcal{C}_{\textup{pr}})$ and $(\mathcal{C}_{\textup{pr}}^{-1},\mathcal{C}_{\textup{pos}}^{-1})$. The prior-preconditioned Hessian $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}$ has been found to be the central object of study in the reduction of finite-dimensional linear Gaussian inverse problems, see [31, 10]. We observe that this operator is directly related to $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ via the equivalent characterisations given by Lemma 3.4 items (i), (ii), (iii) and (iv) and hence to the function $R(\cdot\|\cdot)$ which quantifies the similarity of Gaussian measures by Theorem 3.1(iii).
@@ 3.2 Divergences between equivalent Gaussian measures
To measure the quality of an approximation $\tilde{\mathcal{C}}\in\mathcal{E}$ of $\mathcal{C}_{\textup{pos}}$ and $\tilde{m}_{\textup{pos}}$ of $m_{\textup{pos}}$, we shall use the Rényi divergences of order $\rho\in(0,1)$ and the forward and reverse Kullback–Leibler (KL) divergences. The KL divergence from a measure $\nu_{1}$ to a measure $\nu_{2}$ equivalent to $\nu_{1}$ is defined as
@@
If $\nu_{2}$ is a given measure that needs to be approximated and $\nu_{1}$ is an approximation of $\nu_{2}$, then we refer to $D_{\textup{KL}}(\nu_{2}\|\nu_{1})$ and to $D_{\textup{KL}}(\nu_{1}\|\nu_{2})$ as the ‘forward’ and ‘reverse’ KL divergence of the approximation respectively. The Rényi divergence of order $\rho\in(0,1)$ is defined by
$D_{\textup{Ren},\rho}(\nu_{2}\|\nu_{1})=-\frac{1}{\rho(1-\rho)}\log{\int_{\mathcal{H}}\left(\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}\right)^{\rho}\operatorname{d}\!{\nu}_{1}},$
c.f. [23, eq. (130)]. It holds that $D_{\textup{Ren},\rho}(\nu_{1}\|\nu_{2})=D_{\textup{Ren},1-\rho}(\nu_{2}\|\nu_{1})$, because
$\int_{\mathcal{H}}\left(\frac{\operatorname{d}\!{\nu}_{1}}{\operatorname{d}\!{\nu}_{2}}\right)^{\rho}\operatorname{d}\!{\nu}_{2}=\int_{\mathcal{H}}\left(\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}\right)^{1-\rho}\left(\frac{\operatorname{d}\!{\nu}_{1}}{\operatorname{d}\!{\nu}_{2}}\right)^{1-\rho}\left(\frac{\operatorname{d}\!{\nu}_{1}}{\operatorname{d}\!{\nu}_{2}}\right)^{\rho}\operatorname{d}\!{\nu}_{2}=\int_{\mathcal{H}}\left(\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}\right)^{1-\rho}\operatorname{d}\!{\nu}_{1}.$
This is known as the ‘skew symmetry’ of the Rényi divergence, c.f. [33, Proposition 2]. Consequently, there is no need to consider forward Rényi divergences $D_{\textup{Ren},\rho}(\nu_{2}\|\nu_{1})$ and reverse Rényi divergences $D_{\textup{Ren},\rho}(\nu_{1}\|\nu_{2})$ separately.
In the Gaussian case, an explicit representation of these divergences holds, as shown in [23]. For this, we need a generalisation of the determinant to infinite-dimensional Hilbert spaces. Because in infinite dimensions the eigenvalues of a compact operator accumulate at 0, direct extension of the finite-dimensional definition of the determinant as the product of the eigenvalues to infinite dimensions will result in the determinant function being equal to the constant 00. A generalisation of the concept of the determinant for trace-class and Hilbert–Schmidt operators is given by the Fredholm determinant and Hilbert–Carleman determinant respectively. These are defined on respectively trace-class and Hilbert–Schmidt perturbations of the identity, and are indicated by $\det(I+A)$, $A\in L_{1}(\mathcal{H})$, and respectively $\det_{2}{(I+A)}$, $A\in L_{2}(\mathcal{H})$. We refer to [28, Theorem 3.2, Theorem 6.2] or [29, Lemma 3.3, Theorem 9.2]. For $A\in L_{1}(\mathcal{H})$, we have $\det_{2}{(I+A)}=\det(I+A)\exp(-\operatorname{tr}\,\left({A}\right))$, and for $A\in L_{2}(\mathcal{H})$ the determinants are related via $\det_{2}(I+A)=\det(I+(I+A)\exp(-A)).$ By [28, Theorem 4.2, Theorem 6.2] or [29, Theorem 3.7, Theorem 9.2] for each $\mu\in\mathbb{R}$, we have the expression
@@
where $(\lambda_{i})_{i}$ denotes the eigenvalue sequence of $A$. In the case that $\dim{\mathcal{H}}<\infty$ we note that $A-I\in L_{1}(\mathcal{H})$ and $\det(A)=\det(I+(A-I))=\prod_{i}\lambda_{i}$ and thus $\det(\cdot)$ indeed extends the finite-dimensional definition of the determinant. We can now formulate the explicit expression of the KL and Rényi divergences for equivalent Gaussian measures. The result below holds when $\mathcal{H}$ is a separable Hilbert space of finite or infinite dimension.
@@
Let $m_{1},m_{2}\in\mathcal{H}$ and $\mathcal{C}_{1},\mathcal{C}_{2}\in L_{2}(\mathcal{H})_{\mathbb{R}}$ be positive. If $m_{1}-m_{2}\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ and if $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ satisfies property E, then
@@
The limits above show that the Rényi divergence interpolates the forward KL, obtained in the limit $\rho\uparrow 1$, and the reverse KL, obtained in the limit $\rho\downarrow 0$, between Gaussian measures.
Remark 3.9 (Amari $\alpha$-divergences and Rényi divergences).
The family of Amari $\alpha$-divergences, which is defined for all $\alpha\geq 0$, is another family of divergences which interpolates the forward KL at $\alpha=1$ and reverse KL at $\alpha=0$, c.f. [22, eq. (7)]. For $\alpha\in(0,1)$ and $\alpha>1$, the Amari $\alpha$-divergence $D_{\textup{Am},\alpha}(\nu_{2}\|\nu_{1})$ for equivalent measures $\nu_{1}$ and $\nu_{2}$ on $\mathcal{H}$ is defined by
$D_{\textup{Am},\alpha}(\nu_{2}\|\nu_{1})\coloneqq\frac{-1}{\alpha(1-\alpha)}\left(\int_{\mathcal{H}}\left(\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}\right)^{\alpha}\operatorname{d}\!{\nu}_{1}-1\right),$
and for $\alpha\in(0,1)$ it is related to the $\rho$-Rényi divergence in (13b) with $\rho\leftarrow\alpha$ by
$D_{\textup{Ren},\alpha}(\nu_{2}\|\nu_{1})=\frac{-1}{\alpha(1-\alpha)}\log[1-\alpha(1-\alpha)D_{\textup{Am},\alpha}(\nu_{2}\|\nu_{1})],$
that is,
$D_{\textup{Am},\alpha}(\nu_{2}\|\nu_{1})=\frac{-1}{\alpha(1-\alpha)}\left(\exp[-\alpha(1-\alpha)D_{\textup{Ren},\alpha}(\nu_{2}\|\nu_{1})]-1\right).$ (14)
Since $\nu_{1}$ and $\nu_{2}$ are equivalent and $\alpha\in(0,1)$, $(\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}})^{\alpha}>0$ with $\nu_{1}$-measure 1 and hence $1-\alpha(1-\alpha)D_{\textup{Am},\alpha}(\nu_{2}\|\nu_{1})=\int(\tfrac{\mathrm{d}\nu_{2}}{\mathrm{d}\nu_{1}})^{\alpha}\mathrm{d}\nu_{1}$ is strictly positive. It follows that $D_{\textup{Ren},\alpha}(\nu_{2}\|\nu_{1})$ is a strictly increasing function of $D_{\textup{Am},\alpha}(\nu_{2}\|\nu_{1})$. Thus, for every $10<\alpha<$ 1, minimising the $\alpha$-Rényi divergence corresponds to minimising the Amari $\alpha$-divergence, and vice versa.
Remark 3.10 (Hellinger distance).
Let us denote the Hellinger distance between equivalent measures $\nu_{1}$ and $\nu_{2}$ on $\mathcal{H}$ by $D_{\textup{H}}(\nu_{1},\nu_{2})$, i.e.
$D_{\textup{H}}(\nu_{2},\nu_{1})^{2}\coloneqq\int_{\mathcal{H}}\left(1-\sqrt{\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}}\right)^{2}\operatorname{d}\!{\nu}_{1}=2-2\int_{\mathcal{H}}\sqrt{\frac{\operatorname{d}\!{\nu}_{2}}{\operatorname{d}\!{\nu}_{1}}}\operatorname{d}\!{\nu}_{1}.$
It holds that
$D_{\textup{H}}(\nu_{2},\nu_{1})^{2}=2(1-\exp(-D_{\textup{Ren},1/2}(\nu_{2}\|\nu_{1}))),$ (15)
by e.g. [23, eqs. (134)–(135)], and it follows that minimising the Hellinger distance $D_{\textup{H}}(\nu_{2},\nu_{1})$ is equivalent to minimising the Bhattacharyya distance $D_{\textup{Ren},1/2}(\nu_{2}\|\nu_{1})$ and vice versa.